Circuit boards for wireless telephones are etched, in an acid bath, in batches of 100 boards. A sample of ten boards is randomly selected from each lot for inspection. A particular batch contains one defective board; and x is the number of defective boards in the sample. P( x = 0) is ______
Answer #1

Here x follow binomial with n=10 and p=1/100 then we are to find p(x=0)=8c0*(1/100)^0*(99/100)^10=0.0956